OpenStudy (anonymous):

Solve the equation x − 4 = 9

ganeshie8 (ganeshie8):

add 4 both sides

ganeshie8 (ganeshie8):

x − 4 = 9 +4 +4

OpenStudy (anonymous):

i got 8 and 13 whats next?

ganeshie8 (ganeshie8):

x − 4 = 9 +4 +4

ganeshie8 (ganeshie8):

x = 13

OpenStudy (anonymous):

oooh yeah my bad haha thanks

ganeshie8 (ganeshie8):

thats the answer. -4 + 4 = 0 ;p

ganeshie8 (ganeshie8):

np :)